### v1

# Structural Entropy Guided Unsupervised Graph Out-Of-Distribution Detection

Abstract
With the emerging of huge amount of unlabeled data, unsupervised out-of-distribution (OOD) detection is vital for ensuring the reliability of graph neural networks (GNNs) by identifying OOD samples from in-distribution (ID) ones during testing, where encountering novel or unknown data is inevitable. Existing methods often suffer from compromised performance due to redundant information in graph structures, which impairs their ability to effectively differentiate between ID and OOD data. To address this challenge, we propose SEGO, an unsupervised framework that integrates structural entropy into OOD detection regarding graph classification. Specifically, within the architecture of contrastive learning, SEGO introduces an anchor view in the form of coding tree by minimizing structural entropy. The obtained coding tree effectively removes redundant information from graphs while preserving essential structural information, enabling the capture of distinct graph patterns between ID and OOD samples. Furthermore, we present a multi-grained contrastive learning scheme at local, global, and tree levels using triplet views, where coding trees with essential information serve as the anchor view. Extensive experiments on real-world datasets validate the effectiveness of SEGO, demonstrating superior performance over state-of-the-art baselines in OOD detection. Specifically, our method achieves the best performance on 9 out of 10 dataset pairs, with an average improvement of 3.7% on OOD detection datasets, significantly surpassing the best competitor by 10.8% on the FreeSolv/ToxCast dataset pair.
Introduction
Out-of-distribution (OOD) detection (Yang et al. 2024; Wu et al. 2024; Bao et al. 2024) is a crucial task in machine learning that aims to identify whether a given data point deviates significantly from the training distribution, especially for models deployed in real-world applications where encountering novel or unknown data is inevitable. In graph-based data, the challenge of OOD detection is heightened due to the complex structure and relationships inherent in graphs. In this context, a specific OOD detection model is trained on in-distribution (ID) graphs and then predicts a score for each test sample to indicate its ID/OOD status.
Recent advancements (Wang et al. 2024; Guo et al. 2023; Liu et al. 2023; Yuan et al. 2024b) in graph OOD detection and generation have been explored with growing interest. Several studies (Wang et al. 2024; Guo et al. 2023) employ well-trained graph neural networks (GNNs) (Kipf and Welling 2017; Xu et al. 2019) to fine-tune OOD detectors to identify OOD samples. However, these methods require annotated ID data to pre-train GNNs, which limits their applicability in scenarios where labeled data is unavailable. In contrast, other research (Liu et al. 2023) focuses on training OOD-specific GNN models using only ID data, without relying on any labels or OOD data. They employ unsupervised learning techniques such as graph contrastive learning (GCL) to learn discriminative patterns of unlabeled ID data.
(a) ID and OOD graph samples
(b) Before minimization
(c) After minimization
Figure 1: A toy example of ID and OOD graphs and scoring distributions before/after structural entropy minimization.
Despite the progress made in this area, a challenge still remains less explored. Due to the prevalent presence of redundant information in graph structures, current methods struggle to effectively capture and distinguish the essential structure between ID and OOD data. Without mechanisms to extract substantive information, models are susceptible to irrelevant features and structures that can mislead the learning process. Besides, GCL methods commonly adopt arbitrary augmentations, which may unexpectedly perturb both structural and semantic patterns of the graph, introducing undesired OOD samples and converting ID samples into OOD samples. Although methods like GOOD-D (Liu et al. 2023) attempt to mitigate the issue of structural perturbations through perturbation-free data augmentations, they fail to eliminate the interference of irrelevant information.
Structural entropy (Li and Pan 2016) provides a hierarchical abstraction of graphs to measure the complexity of structure. By minimizing structural entropy, the structural uncertainty of the graph is reduced, which aids in capturing essential information and identifying distinct patterns between ID and OOD samples. We argue that the key to OOD detection is eliminating redundant information in graph structure to focus on the most distinctive and effective information. Fig. 1(a) presents a toy example of ID and OOD graph data, where the light blue and pink shaded areas represent the distinctive parts (i.e., essential information) in ID and OOD graph structure, respectively. Capturing these distinctive parts of the graph can better differentiate OOD samples from ID graphs. We also compute the scoring distributions before and after structural entropy minimization on the BZR/COX2 dataset pair (with BZR as ID dataset and COX2 as OOD dataset). As illustrated in Fig. 1(b) and (c), after the minimization, the OOD scores exhibit smaller variance and a decrease in the overlap of scores between OOD and ID samples. The score frequency density plots show that structural entropy minimization effectively removes redundant information in graph samples, preserving the more distinctive parts of the graph, thus enabling the model to detect distributions more effectively.
In this paper, we propose a novel framework, Structural Entropy guided Graph contrastive learning for unsupervised OOD detection, termed SEGO, to address this challenge. Our approach introduces structural information theory to graph OOD detection for the first time, which can provide significant insights for future research in this field. By minimizing structural entropy, our method effectively removes redundant information of the graph while capturing essential structure. This allows the model to focus on substantive information that distinguishes ID data from OOD data, improving detection performance. Specifically, we extract a coding tree from the original graph using structural entropy minimization to obtain redundancy-eliminated structural information. Additionally, we theoretically demonstrate that minimizing our contrastive loss preserves the maximum mutual information associated with the ground-truth labels. Based on this foundation, we propose a multi-grained contrastive learning scheme using triplet views: the basic view of the original graph, the coding tree as the anchor view representing essential information, and a topological view enriched with topological features. Maximum agreement is achieved at local, global, and tree levels, encouraging the model to encode shared information between these views. Extensive experiments on both real-world datasets demonstrate the superiority of SEGO against state-of-the-art (SOTA) baselines. Our method shows an average improvement of 3.7% in OOD detection across 10 datasets, highlighting its effectiveness in capturing the essential information of graph data for OOD detection tasks. The main contributions of this work are as follows:
  • Guided by structural entropy theory, we propose a novel framework for unsupervised graph OOD detection, termed SEGO, which can remove redundant information and capture the essential structure of graphs, significantly improving the model performance.
  • To mitigate the information gap between node and graph embeddings, we employ a multi-grained contrastive learning scheme using triplet views, which includes coding tree as an anchor view and operates at local, global, and tree levels.
  • Extensive experiments validate the effectiveness of SEGO, demonstrating superior performance over SOTA baselines in OOD detection.
Related Work
Graph Out-of-distribution Detection. OOD detection aims to identify OOD samples from ID ones and has gained increasing traction due to its wide application for vision (Sehwag, Chiang, and Mittal 2021; Wu et al. 2021; Liang, Li, and Srikant 2018) and language (Zhou, Liu, and Chen 2021) data. OOD detection on graph data can be broadly divided into two categories: graph-level (Liu et al. 2023) and node-level detection (Yang et al. 2024; Wu et al. 2024; Bao et al. 2024). Lots of existing methods (Zhu et al. 2024b; Li et al. 2024a, b) focus on improving the generalization ability of GNNs for specific downstream tasks like node classification through supervised learning, rather than identifying OOD samples. Compared to the works (Liang, Li, and Srikant 2018; Hendrycks and Gimpel 2017) relying on ground-truth labels, relatively less effort has been devoted to unsupervised graph-level OOD detection, which remains an urgent research problem. In this work, we provide a novel perspective for identifying OOD graphs by focusing on distinctive essential information based on structural entropy.
Graph Contrastive Learning. As an effective graph self-supervised learning paradigm (Liu et al. 2021, 2022), GCL has achieved great success on unsupervised graph representation learning (Velickovic et al. 2020; Sun et al. 2020a; Hassani and Khasahmadi 2020; You et al. 2020; Zhu et al. 2021; Qiu et al. 2020; Zheng et al. 2022a, b; Ding et al. 2022). Typically, GCL methods involve generating diverse graph views through data augmentation techniques and optimizing the mutual agreement between these views to enhance the representation of samples with similar semantic semantics (You et al. 2020; Zhu et al. 2021; Hassani and Khasahmadi 2020; Zheng et al. 2022b; Ding et al. 2022). However, methods that perform augmentation on graph structures may inadvertently introduce undesired OOD samples within ID data, and views that enhance graph features often suffer from containing redundant information. To effectively capture the essential structure of original graphs, this paper introduces a GCL framework guided by structural entropy, innovatively incorporating triplet views and multi-grained contrast.
Structural Entropy. Structural entropy (Li and Pan 2016), an extension of Shannon entropy (Shannon 1948), quantifies system uncertainty by measuring the complexity of graph structures through the coding tree. Structural entropy has been widely applied in various domains (Wu et al. 2022a, b, 2023; Zhu et al. 2023, 2024a; Hou et al. 2024). In our work, we apply structural entropy in a self-supervised manner to capture the most distinctive part of graphs with essential information for unsupervised graph-level OOD detection.
Notations and Preliminaries
Before formulating the research problem, we first provide some necessary notations. Let $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ represent a graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges. The node features are represented by the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d_{f}}$, where $n=|\mathcal{V}|$ is the number of nodes and $d_{f}$ is the feature dimension. The structure information can also be described by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, so a graph can be alternatively represented by $G=(\mathbf{A},\mathbf{X})$. Notation with its description can be found in Appendix A.
Unsupervised Graph-level OOD Detection.
We consider an unlabeled ID dataset $\mathcal{D}^{id}=\{G_{1}^{id},\cdots,G_{N_{1}}^{id}\}$ where graphs are sampled from distribution $\mathbb{P}^{id}$ and an OOD dataset $\mathcal{D}^{ood}=\{G_{1}^{ood},\cdots,G_{N_{2}}^{ood}\}$ where graphs are sampled from a different distribution $\mathbb{P}^{ood}$. Given a graph $G$ from $\mathcal{D}^{id}$ or $\mathcal{D}^{ood}$, our objective is to detect whether $G$ originates from $\mathbb{P}^{id}$ or $\mathbb{P}^{ood}$. Specifically, we aim to learn a model $f(\cdot)$ that assigns an OOD detection score $s=f(G)$ for an input graph $G$, with a higher $s$ indicating a greater probability that $G$ is from $\mathbb{P}^{ood}$. The model $f$ is trained solely on the ID dataset $\mathcal{D}^{id}_{train}\subset\mathcal{D}^{id}$ and evaluated on a test set $\mathcal{D}^{id}_{test}\cup\mathcal{D}^{ood}_{test}$ (note that $\mathcal{D}^{id}_{test}\cap\mathcal{D}^{id}_{train}=\emptyset$, $\mathcal{D}^{id}_{test}\subset\mathcal{D}^{id}$, and $\mathcal{D}^{ood}_{test}\subset\mathcal{D}^{ood}$).
Structural Entropy.
Structural entropy is initially proposed (Li and Pan 2016) to measure the uncertainty of graph structure, revealing the essential structure of a graph. The structural entropy of a given graph ${G}=\left\{\mathcal{V},\mathcal{E},\mathbf{X}\right\}$ on its coding tree $T$ is defined as:
$\mathcal{H}^{T}({G})=-\sum\limits_{v_{\tau}\in T}\frac{g_{v_{\tau}}}{vol(\mathcal{V})}\log\frac{vol(v_{\tau})}{vol(v_{\tau}^{+})},$ (1)
where $v_{\tau}$ is a node in $T$ except for root node and also stands for a subset $\mathcal{V}_{\tau}\in\mathcal{V}$, $g_{v_{\tau}}$ is the number of edges connecting nodes in and outside $\mathcal{V}_{\tau}$, $v_{\tau}^{+}$ is the immediate predecessor of of $v_{\tau}$ and $vol(v_{\tau})$, $vol(v_{\tau}^{+})$ and $vol(\mathcal{V})$ are the sum of degrees of nodes in $v_{\tau}$, $v_{\tau}^{+}$ and $\mathcal{V}$, respectively.
Graph Contrastive Learning.
In the general graph contrastive learning paradigm for graph classification, two augmented graphs are generated using different graph augmentation operators. Subsequently, representations are generated using a GNN encoder, and further mapped into an embedding space by a shared projection head for contrastive learning. A typical graph contrastive loss, InfoNCE (Chen et al. 2020; Zhu et al. 2021), treats the same graph $G_{i}$ in different views $G_{i}^{\alpha}$ and $G_{i}^{\beta}$ as positive pairs and other nodes as negative pairs. The graph contrastive learning loss $\mathcal{L}_{i}$ of graph $G_{i}$ and total loss $\mathcal{L}$ can be formulated as:
$\vspace{-1.0mm}\begin{gathered}\ell(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{i}^{\beta})=-\operatorname{log}\frac{e^{sim(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{i}^{\beta})/\tau}}{\sum_{j=1,{j\neq i}}^{N}e^{sim(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{j}^{\alpha})/\tau}+e^{sim(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{j}^{\beta})/\tau}},\end{gathered}\vspace{-0.0mm}$ (2)
$\vspace{-1.0mm}\mathcal{L}=\frac{1}{2N}\sum_{i=1}^{N}\Big{[}\ell(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{i}^{\beta})+\ell(\mathbf{z}_{i}^{\beta},\mathbf{z}_{i}^{\alpha})\Big{]},\vspace{-0.0mm}$ (3)
where $N$ denotes the batch size, $\tau$ is the temperature coefficient, and $sim(\cdot,\cdot)$ stands for cosine similarity function.
Figure 2: Overview of our proposed SEGO, which employs multi-grained contrast at local, global, and tree levels using triplet views. The coding tree, obtained by minimizing structural entropy of graph, serves as the essential anchor view that eliminates redundant information. As shown in subfigure on the right, message passing and aggregation on graph are under the guidance of coding tree.
Methodology
In this section, we introduce the framework (see Fig. 2), termed SEGO. We first theoretically reveal that the coding tree with minimum structural entropy effectively captures the essential information of the graph. Additionally, we demonstrate that minimizing our contrastive loss preserves the maximum mutual information associated with ground-truth labels. Based on these insights, we propose a multi-grained contrastive learning using triplet views.
Essential View with Redundancy-eliminated Information
Redundancy-eliminated Essential Information.
The key to effectively distinguishing between ID and OOD graphs lies in maximizing the elimination of redundancy while preserving essential information. According to graph information bottleneck theory (GIB) (Wu et al. 2020), retaining important information in a graph view should involve maximizing mutual information (MI) between the output and labels (i.e., $\max I(f(G);y)$) while reducing mutual information between input and output (i.e., $\min I(G;f(G))$). For unsupervised downstream tasks where ground-truth labels are unavailable, the objective of minimizing $I(G;f(G))$ is to generate an essential view that retains sufficient information while reducing uncertainty (i.e., redundant information) as much as possible. This can be expressed as follows:
[…]
where $I(\cdot;\cdot)$ denotes the MI between inputs.
Definition 1.
The anchor view with redundancy-eliminated essential information is supposed to be a distinctive substructure of the given graph.
Now, let $G^{\ast}$ be the target anchor view of graph $G$, the mutual information between $G$ and $G^{\ast}$ can be formulated as:
$I(G^{\ast};G)=\mathcal{H}(G^{\ast})-\mathcal{H}(G^{\ast}|G),$ (5)
where $\mathcal{H}(G^{\ast})$ is the structural entropy of $G^{\ast}$ and $\mathcal{H}(G^{\ast}|G)$ is the conditional entropy of $G^{\ast}$ conditioned on $G$.
Theorem 1.
The information in $G^{\ast}$ is a subset of information in $G$ (i.e., $\mathcal{H}(G^{\ast})\subseteq\mathcal{H}(G)$); thus, we have:
$\vspace{-3.0mm}\mathcal{H}(G^{\ast}|G)=0.$ (6)
The detailed proof of Theorem 6 is shown in Appendix B.1. Here, the mutual information between $G$ and $G^{\ast}$ can be rewritten as:
$I(G^{\ast};G)=\mathcal{H}(G^{\ast}).$ (7)
Accordingly, to acquire the anchor view with essential information, we need to optimize:
$\min\,I(G;f(G))\Rightarrow\min\,\mathcal{H}(G^{\ast}).$ (8)
Thus, we argue that the view obtained by minimizing structural entropy of a given graph represents the redundancy-eliminated information, serving as an anchor view that retains the graph’s distinctive substructure.
Maximum Effective Mutual Information.
With the essential view eliminating redundant information, we also theoretically prove that our SEGO effectively captures the maximum mutual information between the representations obtained from the anchor view and labels by minimizing the contrastive loss in Eq.3. The InfoMax principle has been widely applied in representation learning literature (Bachman, Hjelm, and Buchwalter 2019; Poole et al. 2019; Tschannen et al. 2020; Yuan et al. 2024a). MI quantifies the amount of information obtained about one random variable by observing the other random variable. We first introduce two lemmas, which are proven in Appendix B.2.
Lemma 1.
Given that $f$ is a GNN encoder with learnable parameters and $G^{\ast}$ is the target anchor view of graph $G$. If $I(f(G^{\ast});f(G))$ reaches its maximum, then $I(f(G^{\ast});G)$ will also reach its maximum.
Lemma 2.
Given the anchor view $G^{\ast}$ of graph $G$ and an encoder $f$, we have
$I(f(G^{\ast});G)\leq I(f(G^{\ast});y)+I(G;G^{\ast}|y).$ (9)
Theorem 2.
Optimizing the loss function in Eq.3 is equivalent to maximizing $I(f(G^{\ast});f(G))$, leading to the maximization of $I(f(G^{\ast});y)$.
Theorem 2 (See proof in Appendix B.3) reveals that our anchor view empowers the encoder to acquire enhanced representations through contrastive learning and preserve more information associated with the ground-truth labels, which will boost the performance of downstream tasks.
Redundancy-aware Multi-grained Triplet Contrastive Learning
Instantiation of Triplet Views.
Our core idea is to capture the essential information of the training ID data and reduce the interference from irrelevant and redundant information. This allows us to distinguish OOD samples during inference by leveraging the different essential information between ID and OOD data.
We first treat a given graph $G_{b}=(\mathbf{A},\mathbf{X})$ as a basic view to directly learn from the original input of the ID data. From this basic view, we construct an anchor view of the graph with minimal structural entropy, as defined by Eq.8. According to structural information theory (Li and Pan 2016), the structural entropy of a graph needs to be calculated with the coding tree. Besides the optimal coding tree with minimum structural entropy (i.e., $\min_{\forall T}\{\mathcal{H}^{T}(G)\}$), a fixed-height coding tree is often preferred for its better representing the fixed natural hierarchy commonly found in real-world networks. Therefore, the $k$-dimensional structural entropy of $G$ is defined using coding tree with fixed height $k$:
$\mathcal{H}^{(k)}({G})=\min_{\forall T:\mathrm{Height}(T)=k}\{\mathcal{H}^{T}({G})\}.$ (10)
The total process of generation of a coding tree with fixed height $k$ can be divided into two steps: 1) construction of the full-height binary coding tree and 2) compression of the binary coding tree to height $k$. Given root node $v_{r}$ of the coding tree $T$, all original nodes in graph ${G}=(\mathcal{V},\mathcal{E})$ are treated as leaf nodes. Correspondingly, based on SEP (Wu et al. 2022a), we design two efficient operators, $\mathbf{MERGE}$ and $\mathbf{DROP}$, to construct a coding tree $T$ with minimum structural entropy. This process is detailed in Algorithm 1 in Appendix C. This tree anchor view $T=\arg\min{\mathcal{H}^{T}({G_{b}})}$ effectively removes redundant information from graphs while preserving distinctive essential structural information, enabling the capture of distinct graph patterns between ID and OOD samples.
In contrastive learning, traditional graph augmentations such as edge modification or node dropping (Velickovic et al. 2020; You et al. 2020) would reduce the model’s sensitivity to OOD data. To address this issue, inspired by GOOD-D (Liu et al. 2023), we adopt a perturbation-free graph augmentation strategy to construct a topological (topo.) view $G_{t}=(\mathbf{A},\mathbf{P})$, where $\mathbf{P}$ is a topological matrix formed by concatenating random walk diffusion and Laplacian positional encoding, formally written as $\mathbf{p}_{i}=[\mathbf{p}_{i}^{(rw)}\,||\,\mathbf{p}_{i}^{(lp)}]$. Specifically, the random walk diffusion encoding is given by $\mathbf{p}_{i}^{(rw)}=[{RW}_{ii},{RW}_{ii}^{2},\cdots,{RW}_{ii}^{r}]\in\mathbb{R}^{r}$, where ${RW}=\mathbf{A}\mathbf{D}^{-1}$ is the random walk transition matrix, and $\mathbf{D}$ is the diagonal degree matrix. The Laplacian positional encoding is defined as $\mathbf{p}_{i}^{(lp)}={\Delta}_{ii}$, where ${\Delta}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$, with $\mathbf{I}$ being the identity matrix. This strategy ensures discriminative representations for OOD detection.
These triplet views in SEGO, namely the basic view $G_{b}$, tree anchor view $T$, and topo. view $G_{t}$, collectively integrate multiple levels of information within the graph. Since the coding tree serves as an abstraction of the essential structure on entire graph, maximizing the MI between tree anchor view $T$ and basic view $G_{b}$ (i.e., $I(T;G_{b})$) allows the model to focus on capturing global redundancy-eliminated patterns. However, relying solely on $T$ and $G_{b}$ might result in the overlooking of fine-grained node-level representations, as $I(T;G_{b})$ primarily emphasizes coarse-grained graph-level information. Therefore, MI between basic view $G_{b}$ and topo. view $G_{t}$ (i.e., $I(G_{b};G_{t})$) is also required, which is captured in both node and graph embeddings. The introduction of $G_{t}$ aligns the information from individual nodes with the overall graph structure, addressing the instability of information. The triplet views in our method mitigate the information gap between node and graph embeddings, effectively capturing both coarse- and fine-grained redundancy-eliminated essential information.
Triplet Views Representing Learning.
To effectively extract embeddings from the basic view $G_{b}$ and topo view $G_{t}$, we utilize two parallel and independent GNN encoders (i.e., encoder $f_{b}$ for the basic view and $f_{t}$ for the topo view) for representation learning, following the approach in GOOD-D (Liu et al. 2023). We employ GIN (Xu et al. 2019) as the backbone for its powerful expression ability. Taking $f_{b}$ as an example, the propagation in the $l$-th layer can be expressed as, $\mathbf{h}_{i}^{(b,l)}=\text{MLP}^{(b,l)}\left(\mathbf{h}_{i}^{(b,l-1)}+\sum\nolimits_{v_{j}\in\mathcal{N}(v_{i})}\mathbf{h}_{j}^{(b,l-1)}\right),$ where MLP is a multilayer perceptron network with 2 layers, $\mathbf{h}_{i}^{(b,l)}$ is the interval embedding of node $v_{i}$ at the $l$-th layer of encoder $f_{b}$, $\mathcal{N}(v_{i})$ is the set of first-order neighbors of node $v_{i}$. After getting node embeddings, we employ a readout function to acquire the graph embedding: $\mathbf{h}_{G}^{(b)}=\sum\nolimits_{v_{i}\in\mathcal{V}_{G}}\mathbf{h}_{i}^{(b)}$, where $\mathcal{V}_{G}$ is the node set of $G$.
For the tree anchor view $T$, the encoder is designed to iteratively transfer messages from the bottom to the top. Formally, the $l$-th layer of the encoder can be written as, $\mathbf{x}_{v}^{(l)}=\text{MLP}^{(l)}\left(\sum\nolimits_{u\in\mathcal{C}(v)}\mathbf{x}_{u}^{(l-1)}\right)$, where $\mathbf{x}^{i}_{v}$ is the feature of $v$ in the $i$-th layer of coding tree $T$, $\mathbf{x}^{0}_{v}$ is the input feature of leaf nodes, and $\mathcal{C}(v)$ refers to the children of $v$. Once the features reach the root node, a readout function is applied to obtain the tree-level embedding $\mathbf{z}_{T}$.
Multi-grained Contrastive Learning Objectives.
To capture the multi-grained mutual information between views, we employ a multi-grained contrastive learning scheme that extracts features at three distinct levels: the local-level for fine-grained feature extraction, the global-level for coarse-grained feature extraction, and the tree-level for capturing essential information of the entire graph. To maximize the agreement between node embeddings from different views of the same graph, we first map $\mathbf{h}_{i}^{(b)}$ and $\mathbf{h}_{i}^{(t)}$ into node-space embeddings $\mathbf{z}_{i}^{(b)}$ and $\mathbf{z}_{i}^{(t)}$ using MLP-based projection networks. The local-level contrast focuses on both inter-view and intra-view node relationships, defined as follows:
$\vspace{-1.2mm}\mathcal{L}_{local}=\frac{1}{|\mathcal{B}|}\sum\limits_{G_{j}\in\mathcal{B}}\frac{1}{2|\mathcal{V}_{j}|}\sum\limits_{v_{i}\in\mathcal{V}_{j}}\left[\ell(\mathbf{z}_{i}^{(b)},\mathbf{z}_{i}^{(t)})+\ell(\mathbf{z}_{i}^{(t)},\mathbf{z}_{i}^{(b)})\right],\vspace{-0.0mm}$ (11)
where $\mathcal{B}$ is a training batch containing multiple graph samples, $\mathcal{V}_{j}$ is the node set of graph $G_{j}$, $\ell(\mathbf{z}_{i}^{(t)},\mathbf{z}_{i}^{(b)})$ and $\ell(\mathbf{z}_{i}^{(b)},\mathbf{z}_{i}^{(t)})$ are calculated following Eq.2.
The global-level contrast allows the model to identify coarse-grained patterns that might be overlooked when focusing solely on finer details:
$\vspace{-1.8mm}\mathcal{L}_{global}=\frac{1}{2|\mathcal{B}|}\sum_{G_{i}\in\mathcal{B}}\Big{[}\ell(\mathbf{z}_{G_{i}}^{(b)},\mathbf{z}_{G_{i}}^{(t)})+\ell(\mathbf{z}_{G_{i}}^{(t)},\mathbf{z}_{G_{i}}^{(b)})\Big{]},\vspace{-0.0mm}$ (12)
where $\mathbf{z}_{G}^{(b)}$ and $\mathbf{z}_{G}^{(t)}$ are transformed from $\mathbf{h}_{G}^{(b)}$ and $\mathbf{h}_{G}^{(t)}$ using MLP-based projection networks.
Tree-level contrast operates at a higher level of abstraction, which can be calculated by:
$\vspace{-1.5mm}\mathcal{L}_{tree}=\frac{1}{2|\mathcal{B}|}\sum_{G_{i}\in\mathcal{B}}\Big{[}\ell(\mathbf{z}_{G_{i}}^{(b)},\mathbf{z}_{T_{i}})+\ell(\mathbf{z}_{T_{i}},\mathbf{z}_{G_{i}}^{(b)})\Big{]}.\vspace{-0.0mm}$ (13)
ID dataset BZR PTC-MR AIDS ENZYMES IMDB-M Tox21 FreeSolv BBBP ClinTox Esol A.A. A.R.
OOD dataset COX2 MUTAG DHFR PROTEIN IMDB-B SIDER ToxCast BACE LIPO MUV
PK-LOF 42.22±8.39 51.04±6.04 50.15±3.29 50.47±2.87 48.03±2.53 51.33±1.81 49.16±3.70 53.10±2.07 50.00±2.17 50.82±1.48 49.63 12.9
PK-OCSVM 42.55±8.26 49.71±6.58 50.17±3.30 50.46±2.78 48.07±2.41 51.33±1.81 48.82±3.29 53.05±2.10 50.06±2.19 51.00±1.33 49.52 12.8
PK-iF 51.46±1.62 54.29±4.33 51.10±1.43 51.67±2.69 50.67±2.47 49.87±0.82 52.28±1.87 51.47±1.33 50.81±1.10 50.85±3.51 51.45 11.1
WL-LOF 48.99±6.20 53.31±8.98 50.77±2.87 52.66±2.47 52.28±4.50 51.92±1.58 51.47±4.23 52.80±1.91 51.29±3.40 51.26±1.31 51.68 10.4
WL-OCSVM 49.16±4.51 53.31±7.57 50.98±2.71 51.77±2.21 51.38±2.39 51.08±1.46 50.38±3.81 52.85±2.00 50.77±3.69 50.97±1.65 51.27 11.1
WL-iF 50.24±2.49 51.43±2.02 50.10±0.44 51.17±2.01 51.07±2.25 50.25±0.96 52.60±2.38 50.78±0.75 50.41±2.17 50.61±1.96 50.87 12.4
OCGIN 76.66±4.17 80.38±6.84 86.01±6.59 57.65±2.96 67.93±3.86 46.09±1.66 59.60±4.78 61.21±8.12 49.13±4.13 54.04±5.50 63.87 7.9
GLocalKD 75.75±5.99 70.63±3.54 93.67±1.24 57.18±2.03 78.25±4.35 66.28±0.98 64.82±3.31 73.15±1.26 55.71±3.81 86.83±2.35 72.23 5.1
InfoGraph-iF 63.17±9.74 51.43±5.19 93.10±1.35 60.00±1.83 58.73±1.96 56.28±0.81 56.92±1.69 53.68±2.90 48.51±1.87 54.16±5.14 59.60 8.5
InfoGraph-MD 86.14±6.77 50.79±8.49 69.02±11.67 55.25±3.51 81.38±1.14 59.97±2.06 58.05±5.46 70.49±4.63 48.12±5.72 77.57±1.69 65.68 7.4
GraphCL-iF 60.00±3.81 50.86±4.30 92.90±1.21 61.33±2.27 59.67±1.65 56.81±0.97 55.55±2.71 59.41±3.58 47.84±0.92 62.12±4.01 60.65 8.7
GraphCL-MD 83.64±6.00 73.03±2.38 93.75±2.13 52.87±6.11 79.09±2.73 58.30±1.52 60.31±5.24 75.72±1.54 51.58±3.64 78.73±1.40 70.70 5.3
GOOD-Dsimp 93.00±3.20 78.43±2.67 98.91±0.41 61.89±2.51 79.71±1.19 65.30±1.27 70.48±2.75 81.56±1.97 66.13±2.98 91.39±0.46 78.68 3.2
GOOD-D 94.99±2.25 81.21±2.65 99.07±0.40 61.84±1.94 79.94±1.09 66.50±1.35 80.13±3.43 82.91±2.58 69.18±3.61 91.52±0.70 80.73 2.2
SEGO 96.66±0.91 85.02±0.94 99.48±0.11 64.42±4.95 80.27±0.92 66.67±0.82 90.95±1.93 87.55±0.13 78.99±2.81 94.59±0.94 84.46 1.1
[…]
During the training phase, we introduce the standard deviation of prediction errors to adaptively adjust the balance of local and global information. This strategy automatically allocates the weights for loss and score terms. Concretely, the overall loss is calculated by:
$\mathcal{L}=\mathcal{L}_{tree}+\sigma_{l}^{\theta}\mathcal{L}_{local}+\sigma_{g}^{\theta}\mathcal{L}_{global},\vspace{-0.0mm}$ (14)
where $\sigma_{l}$ and $\sigma_{g}$ are the standard deviations of predicted errors of the node and graph levels, respectively, and $\theta\geq 0$ is a hyper-parameter that controls the strength of self-adaptiveness, penalizing the term with a larger deviation. During the inference phase, to balance the scores of different levels, we employ z-score normalization based on the mean values and standard deviations of the predicted errors of training samples: $s_{G_{i}}=\frac{s_{l}-\mu_{l}}{\sigma_{l}}+\frac{s_{g}-\mu_{g}}{\sigma_{g}}$, where $\mu_{l}$ and $\mu_{g}$ represent the mean values of the predicted errors at the corresponding levels for the training samples.
Experiment
In this section, we empirically evaluate the effectiveness of the proposed SEGO.  [^1] The code of SEGO is available at https://github.com/name-is-what/SEGO. In particular, the experiments are unfolded by answering the following research questions:
  • RQ1: How effective is SEGO compared with competitive baselines on identifying OOD graphs?
  • RQ2: How transferable is SEGO to anomaly detection?
  • RQ3: How do our multi-grained contrastive losses affect SEGO’s performance?
  • RQ4: How about the parameter sensitivity of SEGO?
Experimental Setups
Datasets. For OOD detection, we employ 10 pairs of datasets from two mainstream graph data benchmarks (i.e., TUDataset (Morris et al. 2020) and OGB (Hu et al. 2020)) following GOOD-D (Liu et al. 2023). We also conduct experiments on anomaly detection settings, where 5 datasets from TUDataset (Morris et al. 2020) are used for evaluation, where the samples in minority class or real anomalous class are viewed as anomalies, while the rest are viewed as normal data. Further details are shown in Appendix D.1.
Baselines. We compare SEGO with 14 competing baseline methods, including 6 GCL (Sun et al. 2020a; You et al. 2020; Liu et al. 2023) based methods, 6 graph kernel based methods (Vishwanathan et al. 2010; Shervashidze et al. 2011), and 2 end-to-end graph anomaly detection methods (Zhao and Akoglu 2021; Ma et al. 2022). More details about implementation are provided in Appendix D.2.
Evaluation and Implementation. We evaluate SEGO with a popular OOD detection metric, i.e., area under receiver operating characteristic Curve (AUC). Higher AUC values indicate better performance. The reported results are the mean performance with standard deviation after 5 runs.
Performance on OOD Detection (RQ1)
To answer RQ1, we compare our proposed methods with 14 competing methods in OOD detection tasks. The AUC results are reported in Table 1. From the comparison results, we observe that SEGO achieves superior performance improvements over the baselines. Specifically, SEGO achieves the best performance on 9 out of 10 dataset pairs and ranks first on average among all baselines with an average rank (A.R.) of 1.1. Additionally, SEGO outperforms all the compared methods in terms of average AUC with a score of 84.46, which is 3.7% higher than the second-best method GOOD-D (Liu et al. 2023). Notably, on the FreeSolv/ToxCast dataset pair, SEGO surpasses the best competitor by 10.8%. While SEGO may not have achieved the absolute best results on the IMDB-M/IMDB-B datasets, its performance is nearly optimal. This is likely because the current coding tree is an approximation that doesn’t fully eliminate all irrelevant redundant information for the high connectivity and edge density of social networks.
Dataset ENZYMES DHFR BZR NCI1 IMDB-B
PK-OCSVM 53.67±2.66 47.91±3.76 46.85±5.31 49.90±1.18 50.75±3.10
PK-iF 51.30±2.01 52.11±3.96 55.32±6.18 50.58±1.38 50.80±3.17
WL-OCSVM 55.24±2.66 50.24±3.13 50.56±5.87 50.63±1.22 54.08±5.19
WL-iF 51.60±3.81 50.29±2.77 52.46±3.30 50.74±1.70 50.20±0.40
GraphCL-iF 53.60±4.88 51.10±2.35 60.24±5.37 49.88±0.53 56.50±4.90
OCGIN 58.75±5.98 49.23±3.05 65.91±1.47 71.98±1.21 60.19±8.90
GLocalKD 61.39±8.81 56.71±3.57 69.42±7.78 68.48±2.39 52.09±3.41
GOOD-Dsimp 61.23±4.58 62.71±3.38 74.48±4.91 59.56±1.62 65.49±1.06
GOOD-D 63.90±3.69 62.67±3.11 75.16±5.15 61.12±2.21 65.88±0.75
SEGO 76.62±7.35 65.31±2.98 89.21±4.51 70.34±1.31 66.48±0.38
[…]
$\mathcal{L}_{tree}$ $\mathcal{L}_{global}$ $\mathcal{L}_{local}$ BZR PTC-MR AIDS ENZYMES IMDB-M Tox21 FreeSolv BBBP ClinTox Esol
COX2 MUTAG DHFR PROTEIN IMDB-B SIDER ToxCast BACE LIPO MUV
- - 54.79±4.08 58.20±3.87 43.68±7.36 49.26±1.11 49.56±5.76 49.26±5.10 49.89±2.95 50.53±0.63 51.97±4.58 54.49±3.57
- - 87.44±4.66 77.84±3.71 97.60±1.05 56.74±1.96 75.22±1.91 65.07±1.32 78.40±6.44 77.66±2.29 70.11±2.44 89.57±2.80
- - 83.51±4.14 72.48±3.77 96.84±0.58 60.85±2.95 79.34±1.81 62.58±0.67 59.48±2.20 69.53±2.29 53.29±4.32 86.49±1.20
- 87.27±8.21 87.71±1.35 97.97±0.04 54.82±2.74 74.51±1.52 64.84±0.29 89.34±0.06 88.34±1.64 79.21±4.55 94.13±1.32
- 79.36±8.69 55.08±1.29 90.66±3.40 63.38±4.18 72.96±3.73 55.68±2.67 61.01±5.29 70.13±0.26 52.14±2.58 77.78±1.01
- 86.29±1.09 77.53±4.03 98.23±0.19 61.55±1.47 75.27±0.54 65.44±1.14 88.04±1.15 80.43±2.58 65.89±4.58 90.94±1.17
96.66±0.91 85.02±0.94 99.48±0.11 64.42±4.95 80.27±0.92 66.67±0.82 90.95±1.93 87.55±0.13 78.99±2.81 94.59±0.94
[…]
Performance on Anomaly Detection (RQ2)
To investigate if SEGO can generalize to the anomaly detection setting (Zhao and Akoglu 2021; Ma et al. 2022), we conduct experiments on 5 datasets following the benchmark in GLocalKD and GOOD-D (Ma et al. 2022; Liu et al. 2023), where only normal data are used for model training. The results are shown in Table $2$More experiments on other 5 datasets for anomaly detection can be found in Appendix D.3. From the results, we find that SEGO shows significant performance improvements compared to other baseline methods. Capturing common patterns in the anomaly detection setting is crucial, which is directly reflected in the performance. Thus, we can conclude that SEGO indeed has a strong capability to learn the essential information of normal graph data.
Ablation Study (RQ3)
Ablation of Multi-grained Contrastive Loss. To address RQ3, we conducted ablation experiments on the OOD detection task by separately removing the different levels of contrastive losses, namely node-, graph-, and tree-level losses. The results are presented in Table 3. Firstly, we observe that applying contrastive loss across all three levels (the last row) achieves the best results on 7 out of 10 datasets and shows promising performance on the remaining datasets. This further elucidates that SEGO better captures the essential information, leading to superior performance in most OOD detection scenarios. Notably, we notice that removing the local-level contrast $\mathcal{L}_{local}$ improves performance on certain dataset pairs (e.g., PTC-MR/MUTAG, BBBP/BACE, and ClinTox/LIPO). These datasets primarily consist of biomolecular data, where molecular activity and interactions are more dependent on the overall topological structure rather than individual node features. Removing $\mathcal{L}_{local}$, which focuses on node-level feature extraction and optimization, reduces interference and allows the model to focus more on the global graph structure, enhancing performance on these datasets.
Effectiveness of MI in Triplet Views. SEGO utilizes $I(T;G_{b})$ at the tree-level contrastive loss. Here, we explore the effectiveness of MI between the anchor view $T$ and topo. view ($I(T;G_{t})$), as well as both views ($I(T;G_{b})\cup I(T;G_{t})$) in identifying OOD graphs. As shown in Fig. 3, we observe that using $I(T;G_{b})$ is more effective in eliminating redundancy from the original graph, whereas, in the view $G_{t}$, data augmentation reintroduces redundant information, leading to sub-optimal performance.
Visualization. We also visualize the embeddings on PTC-MR/MUTAG dataset pair learned by SEGO in triplet views via t-SNE (Van der Maaten and Hinton 2008). As shown in Fig. 4(a)-(c), the embeddings of ID and OOD graphs are well-separated across these views. Among them, the representation gap in $\mathbf{z}_{T}$ is the most pronounced, highlighting the effectiveness of coding tree in extracting essential structures.
Parameter Study (RQ4)
Figure 3: The effectiveness of different views.
(a) w.r.t $\mathbf{z}_{G}^{(b)}$
(b) w.r.t $\mathbf{z}_{G}^{(t)}$
(c) w.r.t $\mathbf{z}_{T}$
Figure 4: T-SNE visualization of embeddings.
The Height $k$ of Graph’s Natural Hierarchy. In the experimental setup, the height $k$ of the coding tree is consistently set to 5, in alignment with the GNN encoder. Here, we delve deeper into the effect of the height $k$ on the graph’s natural hierarchy. The specific performance of SEGO under each height $k$, ranging from 2 to 5, on OOD detection is shown in Fig. A.5. We can observe that the optimal height $k$ with the highest accuracy varies among datasets. We also conducted experiments on the impact of the coding tree height on anomaly detection datasets, as detailed in Appendix D.4.
Figure 5: The natural hierarchy of graph.
Figure 6: The sensitivity of self-adaptiveness strength $\theta$.
Self-adaptiveness Strength $\theta$. To analyze the sensitivity of $\theta$ for SEGO, we alter the value of $\theta$ from 00 to $1$. The AUC w.r.t different selections of $\theta$ is plotted in Fig. 6. We can observe that the variation in AUC with changes in $\theta$ is not entirely consistent across different datasets. This aligns with the findings from the ablation study in Section Ablation Study (RQ3), where we noted that the essential information carried by different datasets varies in their dependence on local node information versus global graph information.
Conclusion
In this paper, we make the first attempt to introduce structural information theory into unsupervised OOD detection regarding graph classification. For this task, we propose a novel structural entropy guided graph contrastive learning framework, termed SEGO, that minimizes structural entropy to capture essential graph information while removing redundant information. Our SEGO employs a multi-grained contrastive learning at node, graph, and tree levels with triplet views, including a coding tree with minimum structural entropy as the anchor view. Extensive experiments on real-world datasets validate the effectiveness of SEGO, demonstrating superior performance over state-of-the-art baselines.
Acknowledgments
This work has been supported by the Guangxi Science and Technology Major Project, China (No. AA22067070), NSFC (Grant No. 61932002) and CCSE project (CCSE-2024ZX-09).
Appendix A Notation Table
As an expansion of the notations in our work, we summarize the frequently used notations in Table A.1.
Notations Descriptions
${G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ Graph with the node set $\mathcal{V}$ and edge set $\mathcal{E}$
$\mathcal{V}$ The set of nodes in the graph
$\mathcal{E}$ The set of edges in the graph
$\mathbf{X}$ The feature matrix
$\mathbb{P}^{id}$, $\mathbb{P}^{ood}$ The distribution where the graphs are sampled from
$y\in\{0,1\}$ The label of the graph
$\mathbf{A}$ The adjacency matrix of the graph
$\mathbf{P}$ The topological matrix of the graph
$G_{b}=\mathbf{(A,X)}$ Basic view of the original graph
$G_{t}=\mathbf{(A,P)}$ Topological view of the graph
$T$ Tree anchor view of the graph
$\mathcal{H}$ The structural entropy
$\mathbf{z}_{i}$ Node-space embedding of node $i$
$\mathbf{z}_{G_{i}}$ Graph-space embedding of graph $G_{i}$
$\mathbf{z}_{T}$ Tree-level embedding
$\mathcal{L}_{local}$, $\mathcal{L}_{global}$, $\mathcal{L}_{tree}$ Local-, global-, tree-level loss
$k$ The coding tree height
$\mathbf{\theta}$ Hyper-parameter for self-adaptiveness
Table A.1: The most frequently used notations in this paper.
Appendix B Theorem Proofs
In this section, we provide proofs for two theorems and the related lemmas.
B.1 Proof of Theorem 1
In this section, we present the proof of the statement $\mathcal{H}(G^{\ast}|G)=0$. First, we repeat the property that the target anchor view should own:
Definition 1.
The anchor view with redundancy-eliminated essential information is supposed to be a distinctive substructure of the given graph.
Now, let $G^{\ast}$ be the target anchor view of graph $G$, the MI between $G$ and $G^{\ast}$ (i.e., $I(G^{\ast};G)$) can be formulated as:
$I(G^{\ast};G)=\mathcal{H}(G^{\ast})-\mathcal{H}(G^{\ast}|G),$ (15)
where $\mathcal{H}(G^{\ast})$ is the structural entropy of $G^{\ast}$ and $\mathcal{H}(G^{\ast}|G)$ is the conditional entropy of $G^{\ast}$ conditioned on $G$. Before the proof, as shown in Fig. A.1, the Venn diagram on the left suggests the mutual information (MI) between the original graph and the anchor view in general cases, while the Venn diagram on the right reveals the MI relationship when the anchor view is a substructure of the given original graph (i.e., complies with Definition 1).
Theorem 1.
The information in $G^{\ast}$ is a subset of information in $G$ (i.e., $\mathcal{H}(G^{\ast})\subseteq\mathcal{H}(G)$); thus, we have:
$\mathcal{H}(G^{\ast}|G)=0.\vspace{-4.0mm}$ (16)
Figure A.1: Venn diagram of the mutual information between original graph and anchor view.
Proof.
According to the definition of Shannon entropy (Shannon 1948), i.e.,
$\mathcal{H}(X)=-\sum_{x\in X}\mathcal{P}(x)\log\mathcal{P}(x),$ (17)
we follow the formulation of graph MI in (Sun et al. 2020b) that $X$ is a set of node representations drawn from an empirical probability distribution of graph $G$, we have:
$\mathcal{H}(G^{\ast}|G)=\mathcal{H}(X^{\ast}|X)$ (18)
So the conditional entropy can be written as:
$\mathcal{H}(X^{\ast}|X)=$ $\sum_{x\in X}\mathcal{P}(x)\mathcal{H}(X^{\ast}|X=x)$ (19)
$=$ $-\sum_{x\in X}\mathcal{P}(x)\sum_{x^{\ast}\in X^{\ast}}\mathcal{P}(x^{\ast}|x)\log\mathcal{P}(x^{\ast}|x)$
$=$ $-\sum_{x\in X}\sum_{x^{\ast}\in X^{\ast}}\mathcal{P}(x^{\ast},x)\log\mathcal{P}(x^{\ast}|x)$
$=$ $-\sum_{x^{\ast},x}\mathcal{P}(x^{\ast},x)\log\mathcal{P}(x^{\ast}|x).$
Considering that $G^{\ast}$ complies with the Definition 1, the illustration of the probability distribution of $G^{\ast}$ and $G$ is shown Fig. A.1. Here, let us first discuss that when $x\in X$ and $x\notin X^{\ast}$, we have
$\mathcal{P}(x^{\ast},x)=0.$ (20)
Therefore, we can transform Eq. 19 to:
$\mathcal{H}(X^{\ast}|X)=$ $-\sum_{x^{\ast},x}\mathcal{P}(x^{\ast},x)\log\mathcal{P}(x^{\ast}|x)$ (21)
$=$ $-\sum_{x^{\ast},x^{\ast}}\mathcal{P}(x^{\ast},x^{\ast})\log\mathcal{P}(x^{\ast}|x^{\ast})$
$-\sum_{x^{\ast},x\notin X^{\ast}}\mathcal{P}(x^{\ast},x)\log\mathcal{P}(x^{\ast}|x)$
$=$ $-\sum_{x^{\ast}\in X^{\ast}}\sum_{x^{\ast}\in X^{\ast}}\mathcal{P}(x^{\ast},x^{\ast})\log\mathcal{P}(x^{\ast}|x^{\ast})$
$=$ $-\sum_{x^{\ast}\in X^{\ast}}\mathcal{P}(x^{\ast})\sum_{x^{\ast}\in X^{\ast}}\mathcal{P}(x^{\ast}|x^{\ast})\log\mathcal{P}(x^{\ast}|x^{\ast})$
$=$ $\sum_{x^{\ast}\in X^{\ast}}\mathcal{P}(x^{\ast})\mathcal{H}(X^{\ast}|X^{\ast}=x^{\ast})$
$=$ $\mathcal{H}(X^{\ast}|X^{\ast})$
$=$ $0$
Therefore, given $\forall x\in X$, we have:
$\mathcal{H}(G^{\ast}|G)=\mathcal{H}(X^{\ast}|X)=0.$ (22)
Accordingly, we have:
$I(G^{\ast};G)=\mathcal{H}(G^{\ast}).$ (23)
B.2 Proof of Lemmas
Lemma 1.
Given that $f$ is a GNN encoder with learnable parameters and $G^{\ast}$ is the target anchor view of graph $G$. If $I(f(G^{\ast});f(G))$ reaches its maximum, then $I(f(G^{\ast});G)$ will also reach its maximum.
Proof.
Because $f(G)$ is a function of $G$,
$I(f(G^{\ast});G)$ $=I(f(G^{\ast});f(G);G)+I(f(G^{\ast});G|f(G))$ (24)
$=I(f(G^{\ast});f(G))+I(f(G^{\ast});G|f(G))$
Thus,
$I(f(G^{\ast});f(G))=I(f(G^{\ast});G)-I(f(G^{\ast});G|f(G))$ (25)
While maximizing $I(f(G^{\ast});f(G))$, either $I(f(G^{\ast});G)$ increases or $I(f(G^{\ast});G|f(G))$ decreases. When $I(f(G^{\ast});G|f(G))$ reaches it minimum value of 0, $I(f(G^{\ast});G)$ will definitely increase. Hence, the process of maximizing $I(f(G^{\ast});f(G))$ can lead to the maximization of $I(f(G^{\ast});G)$ as well. ∎
Lemma 2.
Given the anchor view $G^{\ast}$ of graph $G$ and an encoder $f$, we have
$I(f(G^{\ast});G)\leq I(f(G^{\ast});y)+I(G;G^{\ast}|y).$ (26)
Proof.
$I(f(G^{\ast});G)=I(f(G^{\ast});G;y)+I(f(G^{\ast});G|y)$ (27)
$=$ $I(f(G^{\ast});y)-I(f(G^{\ast});y|G)+I(f(G^{\ast});G|y).$
Due to the non-negativity of mutual information,
$I(G;G^{\ast}|y)$ $=I(f(G^{\ast});G;G^{\ast}|y)+I(G;G^{\ast}|y,f(G^{\ast}))$ (28)
$\geq I(f(G^{\ast});G;G^{\ast}|y)\quad(\text{the non-negativity of}I)$
$=I(f(G^{\ast});G|y)\quad\quad$
According to Eq. 27 and 28, we get
$I(f(G^{\ast});G)+I(f(G^{\ast});y|G)$ $=I(f(G^{\ast});y+I(f(G^{\ast});G|y)$ (29)
$\leq I(f(G^{\ast});y)+I(G;G^{\ast}|y).$
Thus,
$I(f(G^{\ast});G)\leq I(f(G^{\ast});y)+I(G;G^{\ast}|y).$ (30)
B.3 Proof of Theorem 2
To begin with, we revisit the graph contrastive learning loss $\mathcal{L}_{i}$ for graph $G_{i}$ and the total loss $\mathcal{L}$, formally expressed as:
$\vspace{-1.0mm}\begin{gathered}\ell(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{i}^{\beta})=-\operatorname{log}\frac{e^{sim(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{i}^{\beta})/\tau}}{\sum_{j=1,{j\neq i}}^{N}e^{sim(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{j}^{\alpha})/\tau}+e^{sim(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{j}^{\beta})/\tau}},\end{gathered}\vspace{-0.0mm}$ (31)
$\vspace{-1.0mm}\mathcal{L}=\frac{1}{2N}\sum_{i=1}^{N}\Big{[}\ell(\mathbf{z}_{i}^{\alpha},\mathbf{z}_{i}^{\beta})+\ell(\mathbf{z}_{i}^{\beta},\mathbf{z}_{i}^{\alpha})\Big{]},\vspace{-0.0mm}$ (32)
where $N$ denotes the batch size, $\tau$ is the temperature coefficient, and $sim(\cdot,\cdot)$ stands for cosine similarity function.
Theorem 2.
Optimizing the loss function in Eq.32 is equivalent to maximizing $I(f(G^{\ast});f(G))$, leading to the maximization of $I(f(G^{\ast});y)$.
Proof.
Minimizing the contrastive loss in Eq. 32 is equivalent to maximizing a lower bound of the mutual information between the latent representations of two views of the graph, and can be viewed as one way of mutual information maximization between the latent representations (i.e., $\max I(f(G^{\ast});f(G))$). Consequently, the optimization of the loss function in Eq. 32 is equivalent to maximizing $I(f(G^{\ast});f(G))$.
According to Lemma 1, we know that maximizing $I(f(G^{\ast});f(G))$ is equivalent to maximizing $I(f(G^{\ast});G)$, i.e.,
$\max I(f(G^{\ast});f(G))\Rightarrow\max I(f(G^{\ast});G).$ (33)
According to Lemma 2, we have:
$I(f(G^{\ast});G)\leq I(f(G^{\ast});y)+I(G;G^{\ast}|y).$ (34)
Since
$I(G;G^{\ast})=I(G;G^{\ast}|y)+I(G;G^{\ast};y),$ (35)
it follows that
$I(G;G^{\ast}|y)\leq I(G;G^{\ast}).$ (36)
By combining Eq. 34 and Eq. 36, we obtain:
$I(f(G^{\ast});G)$ $\leq I(f(G^{\ast});y)+I(G;G^{\ast}|y)$ (37)
$\leq I(f(G^{\ast});y)+I(G;G^{\ast})$
Thus,
$I(f(G^{\ast});G)-I(G;G^{\ast})\leq I(f(G^{\ast});y).$ (38)
Since in Eq. 23 Theorem 16 we have already proven that $I(G^{\ast};G)=\mathcal{H}(G^{\ast})$, and $G^{\ast}$ is the anchor view of graph $G$ obtained by minimizing structural entropy (i.e., $\min\mathcal{H}(G^{\ast})$), it follows that
$\min I(G;G^{\ast})=\min\mathcal{H}(G^{\ast}).$ (39)
Thus,
$\max I(f(G^{\ast});G)-I(G^{\ast};G)\leq\max I(f(G^{\ast});y).$ (40)
Therefore, minimizing the contrastive loss in Eq. 32 is equivalent to maximizing $I(f(G^{\ast});f(G))$. At this point, $I(f(G^{\ast});G)$ reaches its maximum, and $I(G^{\ast};G)$ reaches its minimum, thereby maximizing $I(f(G^{\ast});y)$.
Appendix C Algorithms
To start with, we initialize a coding tree $T$ by treating all nodes in $\mathcal{V}$ as children of root node $v_{r}$. The construction of the coding tree involves initially building a full-height binary tree with all nodes as leaves and then optimizing this binary tree into a fixed-height coding tree. Specifically, during step 1, an iterative $\mathbf{MERGE}(v_{c}^{1},v_{c}^{2})$ is performed with the goal of minimizing structural entropy to obtain a binary coding tree without height limitation. In this way, selected leaf nodes are combined to form new community divisions with minimal structural entropy. Then, to compress height to a specific hyper-parameter $k$, $\mathbf{DROP}(v_{m})$ is leveraged to merge small divisions into larger ones, and thus the height of the coding tree is reduced, which is still following the structural entropy minimization strategy. Eventually, the coding tree with fixed height $k$ and minimal structural entropy is obtained.
Definition 2.
Assuming $v_{a}$ and $v_{b}$ as two child nodes of root node $v_{r}$, the function $\mathbf{MERGE}(v_{a},v_{b})$ is defined as adding a new node $v_{j}$ as the child of $v_{r}$ and the parent of $v_{a}$ and $v_{b}$:
$\begin{gathered}v_{j}.children=\{v_{a},v_{b}\},\\ v_{r}.children=\{v_{j}\}\cup v_{r}.children.\end{gathered}$ (41)
Since merging nodes in the original graph via the operator $\mathbf{MERGE}(v_{a},v_{b})$ operator reduces the structural entropy of graph $G$, the pair of child nodes to be merged should be the one that maximizes the reduction in structural entropy, formally written as:
$(v_{a},v_{b})=argmax\{\mathcal{H}^{T}({G})-\mathcal{H}^{T_{ab}}({G})|v_{a},v_{b}\in v_{r}.children\}$ (42)
An overview of the $\mathbf{MERGE}$ operator is shown in Fig. A.3. To provide a more detailed explanation of the coding tree construction process, we first revisit the definition of structural entropy as follows:
$\mathcal{H}^{T}({G})=-\sum\limits_{v_{\tau}\in T}\frac{g_{v_{\tau}}}{vol(\mathcal{V})}\log\frac{vol(v_{\tau})}{vol(v_{\tau}^{+})},$ (43)
where $v_{\tau}$ is a node in $T$ except for root node and also stands for a subset $\mathcal{V}_{\tau}\in\mathcal{V}$, $g_{v_{\tau}}$ is the number of edges connecting nodes in and outside $\mathcal{V}_{\tau}$, $v_{\tau}^{+}$ is the immediate predecessor of $v_{\tau}$ and $vol(v_{\tau})$, $vol(v_{\tau}^{+})$ and $vol(\mathcal{V})$ are the sum of degrees of nodes in $v_{\tau}$, $v_{\tau}^{+}$ and $\mathcal{V}$, respectively. When two nodes are merged into a new node (e.g., merging the pair $(v_{a},v_{b})$ into $v_{j}$), the structural information of the new node must satisfy the following equations:
$\begin{gathered}g_{v_{j}}=g_{v_{a}}+g_{v_{b}}-2{Cut}(v_{a},v_{b}),\\ vol(v_{j})=vol(v_{a})+vol(v_{b}),\end{gathered}$ (44)
where ${Cut}(v_{a},v_{b})$ denotes the number of edges cut between nodes when merging $(v_{a},v_{b})$. Fig. A.2 illustrates the process of merging nodes in a graph and adding nodes to the corresponding coding tree through an example.
Definition 3.
Given node $v_{m}$ and its parent node $v_{m}^{+}$ in $T$, the operator $\mathbf{DROP}(v_{m})$ is defined as adding the children of $v_{m}$ and itself to the child set of $v_{m}^{+}$:
$v_{m}^{+}.children=v_{m}^{+}.children\cup v_{m}.children.$ (45)
Similarly, since creating a new node $v_{m}$ through the $\mathbf{DROP}(v_{m})$ operator also changes the structural entropy of graph $G$, the selection of the new node $v_{m}$ should aim to minimize the change in structural entropy, written as:
$v_{m}=argmin\{\mathcal{H}^{T_{m}}({G})-\mathcal{H}^{T}({G})|v_{m}\in T,v_{m}\neq v_{r},v_{m}\notin\mathcal{V}\}$ (46)
An overview of the $\mathbf{DROP}$ operator is shown in Fig. A.4. The construction of coding tree with a fixed height $k$ primarily involves iterations through the two operators to obtain the minimum structural entropy, which is shown in Algorithm 1.
Figure A.2: The process of adding nodes to the corresponding coding tree with an example.
Input: A undirected graph ${G}=(\mathcal{V},\mathcal{E})$ and a specific height $k>1$.
Output: Coding tree ${T}$ with height $k$.
1 Initialize a coding tree $T$ with root node $v_{r}$ and nodes in $\mathcal{V}$ as its children ;
2 // Step 1: full-height binary coding tree construction ;
3 while $|v_{j}.children|>2$ do
4       select child node pair ($v_{a}$,$v_{b}$) using Eq. 42 ;
5       $\mathbf{MERGE}(v_{a},v_{b})$;
6
7 end while
8// Step 2: binary coding tree squeeze to height $k$ ;
9 while $Height(T)>k$ do
10       select node $v_{m}$ using Eq. 46 ;
11       $\mathbf{DROP}(v_{m})$ ;
12
13 end while
14return coding tree $T$ ;
Algorithm 1 Coding tree construction with height $k$ via structural entropy minimization
Figure A.3: Overview of $\mathbf{MERGE}$ to construct full-height binary coding tree.
Figure A.4: Overview of $\mathbf{DROP}$ to squeeze the height of coding tree to $k$.
Given a graph $G=(\mathcal{V},\mathcal{E})$, $|\mathcal{V}|=n$ and $|\mathcal{E}|=m$, the runtime complexity of Algorithm 1 is $O(h_{max}(m\log n+n))$, in which $h_{max}$ is the height of coding tree $T$ after the first stage. In general, the coding tree $T$ tends to be balanced in the process of structural entropy minimization, thus, $h_{max}$ will be around $\log n$. Furthermore, a graph generally has more edges than nodes, i.e., $m\gg n$, thus the runtime of Algorithm 1 almost scales linearly in the number of edges.
Appendix D Experiment
D.1 Dataset Description
Dataset Pair Domain #ID train #ID test #OOD test
BZR / COX2 Molecules 364 41 41
PTC-MR / MUTAG Molecules 309 35 35
AIDS / DHFR Molecules 1800 200 200
Tox21 / SIDER Molecules 7047 784 784
FreeSolv / ToxCast Molecules 577 65 65
BBBP / BACE Molecules 1835 204 204
ClinTox / LIPO Molecules 1329 148 148
Esol / MUV Molecules 1015 113 113
ENZYMES / PROTEINS Proteins 540 60 60
IMDB-M / IMDB-B Social Networks 1350 150 150
Table A.2: Statistics of OOD detection datasets.
Dataset Pair Domain #ID train #ID test #OOD test
BZR Molecules 69 17 64
AIDS Molecules 1280 320 80
COX2 Molecules 81 21 73
NCI1 Molecules 1646 411 411
DHFR Molecules 368 93 59
ENZYMES Proteins 400 100 20
PROTEINS Proteins 360 90 133
DD Proteins 390 97 139
IMDB-B Social Networks 400 100 100
REDDIT-B Social Networks 800 200 200
Table A.3: Statistics of anomaly detection datasets.
For OOD detection, we employ 10 pairs of datasets from two mainstream graph data benchmarks (i.e., TUDataset (Morris et al. 2020) and OGB (Hu et al. 2020)) following GOOD-D (Liu et al. 2023). Specifically, we select 8 pairs of molecular datasets, 1 pair of protein datasets, and 1 pair of social network datasets. $90\%$ of ID samples are used for training, and $10\%$ of ID samples and the same number of OOD samples are integrated together for testing. The partitioning of ID samples for training, along with the division of ID and OOD samples for testing, follows GOOD-D (Liu et al. 2023). Detailed statistics of OOD detection datasets are compiled in Table A.2. Further detailed information about these datasets is categorized and described as follows.
Molecular Datasets
  • BZR (Morris et al. 2020) is a dataset focused on benzodiazepine receptor ligands, containing molecular structures and associated binding affinities. It is crucial for drug design and discovery, specifically for studying receptor-ligand interactions.
  • PTC-MR (Morris et al. 2020) reports the carcinogenicity of 344 chemical compounds in male and female rats and includes 19 discrete labels. It is utilized for predicting the carcinogenic potential of chemical substances.
  • AIDS (Morris et al. 2020) contains data on anti-HIV compounds, including their molecular structures and biological activities, serving as a valuable resource for the development of anti-HIV drugs.
  • ENZYMES (Morris et al. 2020) is a dataset consisting of protein structures classified into enzyme types based on their functionality. It is used for protein function prediction and enzyme classification.
  • COX2 (Morris et al. 2020) comprises data on cyclooxygenase-2 inhibitors, which are compounds with anti-inflammatory properties. This dataset is essential for research and development of anti-inflammatory drugs.
  • MUTAG (Morris et al. 2020) has seven kinds of graphs derived from 188 mutagenic aromatic and heteroaromatic nitro compounds. It is used for studying the mutagenicity of chemical substances.
  • DHFR (Morris et al. 2020) The DHFR dataset includes dihydrofolate reductase inhibitors, important in the development of antibacterial and anticancer drugs, aiding in drug discovery and medicinal chemistry research.
  • PROTEINS (Morris et al. 2020) contains data on protein structures and their functionalities. Nodes represent secondary structure elements (SSEs), and edges connect neighboring elements in the amino acid sequence or 3D space. This dataset is used for protein structure prediction and functional analysis.
  • Tox21 (Hu et al. 2020) is a dataset containing toxicity data on 12 biological targets, which has been used in the 2014 Tox21 Data Challenge and includes nuclear receptors and stress response pathways.
  • BBBP (Hu et al. 2020; Martins et al. 2012) includes records of whether a compound has the permeability property of penetrating the blood-brain barrier, essential for the design of central nervous system drugs.
  • ClinTox (Hu et al. 2020; Novick et al. 2013; Gayvert, Madhukar, and Elemento 2016) contains clinical toxicity data on a variety of drug compounds, classifying drugs approved by the FDA and those that have failed clinical trials for toxicity reasons.
  • ToxCast (Hu et al. 2020; Richard et al. 2016) includes high-throughput screening data on the toxicity of chemical substances, with measurements based on over 600 in vitro screenings. This dataset is used for large-scale toxicity assessment and environmental health research.
  • SIDER (Hu et al. 2020; Kuhn et al. 2016) contains information on drug side effects, grouped into 27 system organ classes, also known as the Side Effect Resource. It is utilized for predicting drug side effects and improving drug safety profiles.
  • BACE (Hu et al. 2020; Subramanian et al. 2016) includes qualitative binding results for a set of inhibitors of human $\beta$-secretase 1, which are potential treatments for Alzheimer’s disease. This dataset is used in Alzheimer’s disease research and drug development.
  • FreeSolv (Hu et al. 2020) includes data on the hydration free energy of small molecules, used for molecular dynamics simulations and solubility studies.
  • Esol (Hu et al. 2020) contains data on the aqueous solubility of compounds, used for studying compound solubility and drug design.
  • LIPO (Hu et al. 2020) includes data on the lipophilicity of chemical compounds. It is used for studying the partitioning of compounds between water and oil phases, which is important in drug design.
  • MUV (Hu et al. 2020; Gardiner et al. 2011) includes data on the activity of compounds from virtual screening, designed for validation of virtual screening techniques.
  • HIV (Hu et al. 2020) contains experimentally measured abilities to inhibit HIV replication.
Protein Datasets
  • PROTEINS (Morris et al. 2020) contains data on protein structures and their functionalities. Nodes represent secondary structure elements (SSEs), and edges connect neighboring elements in the amino acid sequence or 3D space. This dataset is used for protein structure prediction and functional analysis.
  • ENZYMES (Morris et al. 2020) is a dataset consisting of protein structures classified into enzyme types based on their functionality. It is used for protein function prediction and enzyme classification.
Social Network Datasets
  • IMDB-BINARY (Morris et al. 2020) (abbreviated as IMDB-B) is derived from the collaboration of a movie set. Each graph consists of actors or actresses, with edges representing their cooperation in a movie. The label corresponds to movie’s genre. This dataset is used for movie classification and recommendation system studies.
  • IMDB-MULTI (Morris et al. 2020) (abbreviated as IMDB-M) consists of graphs derived from movie collaborations which is similar to IMDB-BINARY, but with multi-class labels. It is utilized in multi-class classification tasks in social network analysis.
We also conduct experiments on anomaly detection (AD) settings, where 10 datasets from TUDataset (Morris et al. 2020) are used for evaluation. Following the setting in GlocalKD (Ma et al. 2022), the samples in the minority class or real anomalous class are viewed as anomalies, while the rest are viewed as normal data. Similar to (Ma et al. 2022; Zhao and Akoglu 2021), only normal data are used for model training. Detailed statistics of anomaly detection datasets are compiled in Table A.3.
ID Dataset PROTEINS-full ENZYMES AIDS DHFR BZR COX2 DD NCI1 IMDB-B REDDIT-B A.A A.R.
PK-OCSVM 50.49±4.92 53.67±2.66 50.79±4.30 47.91±3.76 46.85±5.31 50.27±7.91 48.30±3.98 49.90±1.18 50.75±3.10 45.68±2.24 49.46 9.14
PK-iF 60.70±2.55 51.30±2.01 51.84±2.87 52.11±3.96 55.32±6.18 50.05±2.06 71.32±2.41 50.58±1.38 50.80±3.17 46.72±3.42 54.07 7.73
WL-OCSVM 51.35±4.35 55.24±2.66 50.12±3.43 50.24±3.13 50.56±5.87 49.86±7.43 47.99±4.09 50.63±1.22 54.08±5.19 49.31±2.33 50.94 8.27
WL-iF 61.36±2.54 51.60±3.81 61.13±0.71 50.29±2.77 52.46±3.30 50.27±0.34 70.31±1.09 50.74±1.70 50.20±0.40 48.26±0.32 54.66 7.50
GraphCL-iF 60.18±2.53 53.60±4.88 79.72±3.98 51.10±2.35 60.24±5.37 52.01±3.17 59.32±3.92 49.88±0.53 56.50±4.90 71.80±4.38 59.43 6.73
OCGIN 70.89±2.44 58.75±5.98 78.16±3.05 49.23±3.05 65.91±1.47 53.58±5.05 72.27±1.83 71.98±1.21 60.19±8.90 75.93±8.65 65.69 4.91
GLocalKD 77.30±5.15 61.39±8.81 93.27±4.19 56.71±3.57 69.42±7.78 59.37±12.67 80.12±5.24 68.48±2.39 52.09±3.41 77.85±2.62 69.60 3.55
GOOD-Dsimp 74.74±2.28 61.23±4.58 94.09±1.75 62.71±3.38 74.48±4.91 60.46±12.34 72.24±1.82 59.56±1.62 65.49±1.06 87.87±1.38 71.29 3.36
GOOD-D 71.97±3.86 63.90±3.69 97.28±0.69 62.67±3.11 75.16±5.15 62.65±8.14 73.25±3.19 61.12±2.21 65.88±0.75 88.67±1.24 72.26 2.27
SEGO 75.71±3.95 76.62±7.35 96.46±1.55 65.31±2.98 89.21±4.51 60.93±5.01 73.61±2.22 70.34±1.31 66.48±0.38 88.11±2.13 76.28 1.55
[…]
D.2 Baselines
We evaluate the performance of SEGO by comparing it against 14 state-of-the-art baseline methods. A detailed discussion of these methods is provided below.
GCL+detector. This category of methods first generates representations using state-of-the-art (SOTA) GCL techniques and then distinguishes OOD samples using detectors based on these learned representations. We utilize two graph-level GCL methods, InfoGraph (Sun et al. 2020a) and GraphCL (You et al. 2020), for representation learning. Alongside the iF detector (Liu, Ting, and Zhou 2008), we also incorporate the Mahalanobis distance-based (MD) detector, which has demonstrated effectiveness in detecting OOD data (Sehwag, Chiang, and Mittal 2021; Zhou, Liu, and Chen 2021). GOOD-D (Liu et al. 2023), a GCL-based method that has shown strong performance in unsupervised OOD detection tasks, is included as the primary SOTA method in our comparison.
Graph kernel+detector. This type of methods initially extract vectorized representations using graph kernels (Vishwanathan et al. 2010) and then employ OOD/anomaly detectors to identify OOD samples based on these representations. We use the Weisfeiler-Lehman kernel (WL) (Shervashidze et al. 2011) and the propagation kernel (PK) (Neumann et al. 2016) for the kernel-based representations, and we apply local outlier factor (LOF) (Breunig et al. 2000), one-class SVM (OCSVM) (Manevitz and Yousef 2001), and isolation forest (iF) (Liu, Ting, and Zhou 2008) as detectors.
End-to-end. We also compare our approach with two graph anomaly detection methods that are trained in an end-to-end manner. The first method is OCGIN (Zhao and Akoglu 2021), which optimizes a GIN encoder using an SVDD objective. The second method is GLocalKD (Ma et al. 2022), which identifies anomalies through knowledge distillation.
For the experiment environment, we use the following software framework: Python 3.7, Pytorch 1.8, CUDA 11.0, and Pytorch-Geometric 2.0.1. The hardware setup includes an Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz, 256GB RAM, and a Tesla V100 PCIe 32GB GPU.
D.3 Performance on Anomaly Detection
To investigate if SEGO can generalize to the anomaly detection setting (Zhao and Akoglu 2021; Ma et al. 2022), we conduct experiments on 10 datasets following the benchmark in GLocalKD and GOOD-D (Ma et al. 2022; Liu et al. 2023), where only normal data are used for model training. The results are shown in Table A.4. From the results, we find that our method achieves the best performance on 4 out of 10 datasets and runner-up performance on 6 datasets, with an average improvement of 4.02% over the SOTA. This demonstrates that SEGO indeed has a strong capability to transfer to anomaly detection tasks. Thus, we can conclude that capturing common patterns in the anomaly detection setting is crucial, which is directly reflected in the performance.
D.4 Impacts of The Natural Hierarchy of Graph on Anomaly Detection.
We also conducted experiments on the impact of the coding tree height $k$ on 5 anomaly detection datasets, as shown in Fig. A.5. We observe that the impact of coding tree height on anomaly detection performance varies slightly across different datasets. Notably, on the ENZYMES, DHFR, and BZR datasets, the best performance is achieved when the default experimental setting of $k=5$ is used.
Figure A.5: The natural hierarchy of graph on AD datasets.